OpenStudy (anonymous):

simplify- use exact answers only.. sq root of 121

OpenStudy (anonymous):

11

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

and for 80 would it be 6 outside and 8 under the sq rt

OpenStudy (anonymous):

?

OpenStudy (anonymous):

im sorry, i dont get what your asking...

OpenStudy (anonymous):

\[\sqrt{80}=\sqrt{5*16}=4\sqrt{5}\]

OpenStudy (anonymous):

oohhh ^^ well there yu go :D